OpenStudy (anonymous):

Which ordered pair in the form (a, b) is a solution of this equation? 7a – 5b = 28 A. (4, 0) B. (–2, –3) C. (–3, –2) D. (0, 4)

OpenStudy (danjs):

just plug in (a,b)=(4,0) or the other 3 and see which one equals 28

OpenStudy (michele_laino):

please you have to substitute all of your pairs into your equation

OpenStudy (danjs):

7(4) - 5(0) = 28

OpenStudy (anonymous):

Thanks Dan

OpenStudy (jordanloveangel):

try plugging the points, example first point you have (0,4) 7a - 5b = 28 7(0) - 5(4) = 28 0 - 20 = 28 -20 = 28

OpenStudy (jordanloveangel):

so if you replace the numbered pairs with the a and b sooner or later u will get the answer do u get it now